# Homophily-Driven Sanitation View for Robust Graph Contrastive Learning

### v2

@@ VI-C1 Gumbel-Softmax Re-parametrization
VI-C2 Projection Gradient Descent
Formally, to optimize the parameters $\mathbf{P}$ under the constraint $\mathcal{S}_{1}$ defined in Eqn. 19, we deploy the projection gradient descent to optimize the relaxed constrained optimization problem introduced in Eqn. 19:
at the $t$-th iteration, where $\alpha$ is the learning rate of the projection gradient descent, $\nabla_{\mathbf{P}}\mathcal{L}(t_{1}(\mathcal{G}),t_{2}(\mathcal{G}),\theta)$ denotes the gradient of the loss defined in Eqn. 16, $\prod_{\mathcal{S}_{1}}(\cdot)$ is the projection operator to project the updated parameters to satisfy the constraint. Referring to [36], the projection operator $\prod_{\mathcal{S}_{1}}(\cdot)$ has the closed-form solutions:
where $\prod_{[0,1]}[\mathbf{P}]$ clips the parameters vector $\mathbf{P}$ into the range $[0,1]$. The bisection method [37] is used over $\mu\in\{\min(\mathbf{P}-\mathbf{1}_{E}),\max(\mathbf{P})\}$ to find the solution to $\sum_{e}\prod_{[0,1]}[\mathbf{P}-\mu\mathbf{1}_{E}]=\epsilon$ with the convergence rate $\log_{2}[(\max(\mathbf{P})-\min(\mathbf{P}-\mathbf{1}_{E}))/\xi]$ with $\xi$-error tolerance [38].
VI-D Unsupervised Hyperparameter Tuning
One issue that remains unsolved is the tuning of the hyperparameter $\eta$. While some previous works [6, 5, 4] use node labels for tuning hyperparameters, we consider a strict unsupervised setting where node labels are not available during training and are only used in the downstream node classification task for testing. We thus propose a fully unsupervised method to tune the parameter $\eta$.
In Sec. III-C, we introduced the normalized cut loss in Eqn. 2, which counts the number of inter-cluster links within a given graph. The loss value depends on the quality of the clustering result determined by cluster assignment matrix $\mathbf{C}$. We propose to use this normalized cut loss as the supervision for searching the parameter $\eta$. The intuition is that if the generated node embeddings have high quality, the clustering result should be accurate, leading to a low loss value. In fact, several previous works [39, 28] have utilized this loss as supervision in different ways.
Following this intuition, we use the normalized cut loss as a metric to quantify the quality of node embedding and tune $\eta$ by monitoring this loss during training. Specifically, during the $t$-th iteration, we obtain the sanitized adjacency matrix $\mathbf{A}_{\mathcal{S}}^{(t)}$ (corresponding with $\mathbf{E}_{\mathcal{S}}^{(t)}$ in Eqn. (19)). Then we feed it into a two-layered GNN encoder to the sanitized node embeddings $\mathbf{H}_{\mathcal{S}}^{(t)}=f_{\theta}(\mathbf{A}_{\mathcal{S}}^{(t)},\mathbf{X})$ from Eqn. 1 and get the pseudo normalized cut loss:
@@ VI-D Unsupervised Hyperparameter Tuning
$\begin{split}\mathcal{L}_{pnc}^{(t)}=\operatorname{Tr}((\sigma(\mathbf{H}_{1}^{(t)})^{\top}\mathbf{L}^{(t)}\sigma(\mathbf{H}_{1}^{(t)}))\oslash(\sigma(\mathbf{H}_{1}^{(t)})^{\top}\mathbf{D}^{(t)}\sigma(\mathbf{H}_{1}^{(t)})),\end{split}$
where $\sigma(\cdot)$ is the sigmoid function. $\mathcal{L}_{pnc}$ is time-efficient since it does not require additional models or parameters to be trained. We use the early-stopping trick [40] to effectively tune the hyperparameter $\eta$. That is,
Figure 3: $\mathcal{L}_{pnc}^{t^{*}}$ v.s. Accuracy.
@@
we record the value of $\mathcal{L}_{pnc}^{(t)}$ for each iteration during training and select the model with the minimum $\mathcal{L}_{pnc}$ as our final result for evaluation. We denote the best metric as $\mathcal{L}_{pnc}^{t^{*}}$. Fig. 3 presents the values of $\mathcal{L}_{pnc}^{t^{*}}$ and node classification accuracy with varying $\eta$. The results indicate that in most cases $\mathcal{L}_{pnc}^{t^{*}}$ is negatively related to the node classification accuracy, especially when the value of $\mathcal{L}_{pnc}^{t^{*}}$ falls into the neighborhood of the minimum value, i.e., $\eta\in[0.1,3]$ (the Pearson correlation coefficients $r(\mathcal{L}_{pnc}^{t^{*}},accuracy)=-0.89$). This phenomenon demonstrates that it is reasonable to tune $\eta$ via searching for the best $\mathcal{L}_{pnc}^{t^{*}}$, since the GCL model that achieves the minimum $\mathcal{L}_{pnc}^{t^{*}}$ is likely to get the best accuracy. The algorithm of our work is in Alg. 1.
Input: Poisoned graph $\mathcal{G}=\{\mathbf{E},\mathbf{X}\}$, $\eta$.
Parameters: Sanitation probability $\mathbf{P}$, GNN encoder $f_{\theta}(\cdot)$ with weight matrices $\theta=\{\mathbf{W}^{(1)},\mathbf{W}^{(2)}\}$.
Output: Sanitized node embeddings matrix $\mathbf{H}_{1}$.
1:  Let $t=0$, initialize parameters $\mathbf{P}=\mathbf{P}^{0}$, $\theta=\theta^{0}$.
2:  while $t\leq T$ do
3:     Sample $g\sim Gumbel(0,1)$.
4:     Sample sanitation view $\mathcal{G}_{1}^{t}$ from $\mathbf{E}_{\mathcal{S}}^{t}=\mathbf{E}\odot(\mathbf{1}_{E}-\lfloor\frac{1}{1+e^{-(\log\mathbf{P}^{t}+g)/\tau}}+\frac{1}{2}\rceil)$.
5:     Sample another view $\mathcal{G}_{2}^{t}$ by randomly dropping links.
6:     Compute graph Laplacian $\mathbf{L}^{t}=\mathbf{I}-(\tilde{\mathbf{D}}^{t})^{-\frac{1}{2}}\tilde{\mathbf{A}}^{t}(\tilde{\mathbf{D}}^{t})^{-\frac{1}{2}}$.
7:     Compute loss $\mathcal{L}=\mathcal{L}_{info}(\mathcal{G}_{1}^{t},\mathcal{G}_{2}^{t})+\eta\operatorname{Tr}(\mathbf{X}^{\top}\mathbf{L}^{t}\mathbf{X})$.
8:     Projection gradient descent and update $\mathbf{P}^{t}$ from Eqn. 20.
9:     Gradient descent and update $\theta^{t}$ based on $\mathcal{L}$.
10:     Compute pseudo normalized cut loss $\mathcal{L}_{pnc}^{t}$ from Eqn. VI-D.
11:  end while
12:  Get the best iteration $t^{*}$ with the minimum value $\mathcal{L}_{pnc}^{t^{*}}$.
13:  return  Sanitized node embeddings $\mathbf{H}_{1}=f_{\theta^{t^{*}}}(\mathbf{A}^{t^{*}},\mathbf{X})$.
Algorithm 1 GCHS
VII Experiments
@@
In this section, we introduce the experimental settings and conduct comprehensive experiments to validate the effectiveness of our robust model. Specifically, we outline the questions that we seek to answer below.
@@
  • RQ1: Comparing with baseline models, does our model achieve better robust performance for node classification and graph clustering downstream tasks?
@@
VII-A Experimental settings
VII-A1 Datasets
We use six popular benchmark graph datasets: Cora, CiteSeer, Cora-ML, Photo, Computers and WikiCS [10, 41, 42] for evaluation. The detailed statistics of these datasets are summarized in Tab. II.
Table II: Dataset statistics.
@@ VII-A1 Datasets
Cora $2708$ $5278$ $7$ $1433$
Citeseer $3327$ $4552$ $6$ $3703$
Cora-ML $2995$ $8416$ $7$ $2879$
Photo $7650$ $119081$ $8$ $745$
@@ VII-A2 Settings
To validate the quality of the node embeddings effectively, we utilize semi-supervised node classification and unsupervised graph clustering as downstream tasks. For node classification, we train a multi-class logistic regression model with Adam optimizer [43] and randomly split the node labels into three groups: training, validation and testing set with the ratio as $10\%$, $10\%$ and $80\%$. For graph clustering, the K-means clustering [44] algorithm is run based on the pre-trained node embeddings derived from GCL models and we report the normalized mutual information (NMI) to quantify the robustness of the graph clustering performances. A two-layered GNN encoder is deployed for all the GCL models. The embedding dimension of the first layer is $128$ for all the datasets and the embedding dimension of the second layer is $32$ for Cora, CiteSeer and Cora-ML and $128$ for Photo, Computers and WikiCS. We train the GNN encoder parameters $\theta$ via Adam optimizer with the learning rate equal to $0.0005$ and the training epochs equal to $1000$. We run all the models $10$ times with different seeds and report the mean value for a fair comparison. The hyperparameter $\eta$ based on $\mathcal{L}_{pnc}^{*}$ in Eqn. VI-D without querying the node labels. All the models are executed on an Intel 10C20T Core i9-10850K CPU with GIGABYTE RTX3090 24GB GPU.
Table III: Performances on node classification against Mettack.
@@
Dataset $\frac{B}{|E|}$ BGRL DGI MVGRL GRACE ARIEL GCA SPAN GCHS
Cora 0% 75.12(1.7) 81.04(1.7) 82.93(0.8) 81.02(1.17) 81.14(0.87) 82.09(0.36) 81.83(0.88) 80.03(0.78)
5% 66.97(1.4) 73.41(1.0) 75.73(0.5) 72.46(1.80) 74.92(1.23) 73.80(1.09) 74.18(0.94) 78.78(1.02)
10% 60.11(1.7) 68.03(1.8) 71.10(1.4) 65.37(2.80) 71.93(1.24) 70.26(1.34) 68.81(1.89) 75.93(0.85)
15% 51.85(3.0) 59.70(2.7) 64.69(1.9) 52.62(3.10) 52.91(4.39) 57.89(1.52) 58.44(3.23) 71.89(1.56)
20% 44.18(4.3) 51.43(2.3) 57.29(2.8) 41.17(3.98) 43.71(3.87) 46.03(3.01) 45.71(2.50) 68.81(2.44)
CiteSeer 0% 62.55(1.4) 70.63(1.3) 72.35(0.8) 70.97(1.14) 71.72(1.26) 72.37(1.06) 71.61(0.96) 71.80(1.28)
5% 56.90(2.7) 66.43(1.1) 68.93(0.9) 69.55(1.96) 70.79(1.70) 67.93(1.14) 69.08(1.32) 70.93(1.27)
10% 52.30(2.6) 60.78(2.8) 65.01(1.2) 63.49(2.25) 69.14(2.14) 60.82(1.26) 65.18(2.58) 69.43(1.28)
15% 44.44(2.7) 53.30(2.4) 58.51(1.3) 53.67(2.73) 56.46(2.33) 52.52(1.67) 56.08(3.50) 65.84(1.16)
20% 40.89(3.2) 49.54(1.9) 51.69(1.6) 47.62(2.95) 47.93(3.17) 48.79(1.42) 50.68(3.02) 62.65(1.48)
Cora-ML 0% 74.74(1.3) 81.93(0.9) 82.93(0.9) 82.46(0.54) 82.68(0.44) 82.78(0.76) 82.03(0.67) 81.41(0.75)
5% 64.26(1.9) 67.06(2.9) 70.18(2.5) 68.65(1.96) 71.21(1.74) 69.84(1.74) 69.34(0.88) 77.90(0.83)
10% 53.08(2.6) 49.34(2.1) 54.20(2.4) 50.49(1.72) 51.87(1.76) 46.81(1.86) 50.44(1.22) 71.57(1.66)
15% 44.09(3.6) 42.21(1.5) 45.96(3.3) 40.89(3.40) 41.11(2.31) 37.17(3.58) 41.77(2.37) 64.56(1.82)
20% 37.41(3.3) 34.35(0.4) 37.40(2.3) 31.57(3.28) 28.52(2.22) 26.69(2.10) 33.46(2.47) 56.43(2.50)
Photo 0% 85.79(0.4) 78.26(0.5) 89.48(0.8) 90.16(0.2) 89.39(0.7) 89.19(0.9) 87.53(1.0) 90.87(0.3)
5% 66.50(2.8) 62.24(3.4) 73.14(1.3) 70.87(1.1) 69.85(2.4) 76.45(2.1) 72.46(2.3) 78.37(2.2)
10% 52.62(3.9) 47.76(3.7) 58.63(3.3) 51.39(1.5) 53.92(1.8) 60.57(2.0) 63.46(0.4) 67.80(2.4)
15% 43.56(2.4) 40.73(3.6) 49.82(2.0) 44.37(1.6) 45.98(1.8) 51.66(1.4) 50.02(0.4) 59.19(1.5)
20% 35.68(3.5) 37.91(4.0) 45.00(2.0) 32.45(1.4) 34.19(1.8) 45.35(1.3) 46.81(3.4) 54.68(2.9)
Computers 0% 77.32(1.6) 72.88(2.9) 80.03(0.6) 84.30(0.8) 83.83(1.4) 85.02(1.2) 80.87(0.8) 83.99(1.1)
5% 63.12(1.3) 63.91(1.9) 69.16(0.6) 66.22(1.0) 67.51(1.7) 68.02(0.9) 64.06(1.4) 72.25(2.2)
10% 61.33(2.1) 57.06(1.9) 64.58(0.4) 58.07(2.0) 57.04(1.5) 58.90(2.1) 54.81(1.3) 66.14(2.1)
15% 52.59(3.4) 51.40(2.0) 54.99(2.6) 52.58(2.1) 43.65(2.6) 52.90(2.6) 45.63(3.2) 59.93(3.8)
20% 37.61(2.3) 41.54(4.9) 44.98(3.5) 37.21(3.9) 34.71(1.7) 38.81(3.6) 37.39(2.1) 51.13(4.1)
WikiCS 0% 74.78(0.5) 77.09(0.8) 78.42(0.4) 79.70(0.2) 78.48(0.5) 79.35(0.3) 79.27(0.4) 78.02(0.3)
5% 40.91(1.0) 43.71(2.2) 42.17(1.8) 41.30(1.1) 43.92(1.0) 41.67(0.8) 40.60(1.1) 55.93(0.8)
10% 29.93(0.9) 33.88(2.7) 30.73(0.7) 30.89(1.1) 33.55(1.4) 31.18(0.7) 28.48(0.7) 49.30(0.6)
15% 24.35(0.7) 27.10(2.0) 24.54(1.1) 25.40(0.4) 27.37(0.6) 25.25(0.5) 23.32(0.7) 42.96(1.1)
20% 21.60(1.8) 23.79(1.3) 22.38(0.8) 23.23(0.4) 24.56(0.7) 23.03(0.7) 21.79(0.5) 36.47(1.0)
Table IV: Performances on node classification against CLGA.
@@
Dataset $\frac{B}{|E|}$ BGRL DGI MVGRL GRACE ARIEL GCA SPAN GCHS
Cora 5% 70.78(1.1) 78.96(1.3) 80.09(0.9) 78.95(0.93) 78.47(0.72) 78.55(0.72) 78.96(0.72) 79.13(0.59)
10% 68.35(1.9) 75.64(1.2) 77.92(0.9) 76.79(1.35) 76.31(0.92) 77.15(0.78) 77.22(1.09) 78.62(1.11))
15% 66.15(1.3) 73.72(0.7) 75.89(1.0) 75.38(1.14) 74.75(0.89) 74.71(1.15) 76.11(1.19) 78.18(1.02)
20% 64.44(1.8) 71.94(1.7) 74.64(0.8) 73.13(1.06) 74.44(1.19) 71.78(1.00) 74.43(1.64) 76.82(1.16)
CiteSeer 5% 58.52(1.3) 66.24(1.5) 68.60(1.2) 69.38(1.22) 70.43(1.17) 67.92(1.04) 70.38(1.11) 71.18(1.22)
10% 56.27(1.3) 65.62(1.5) 67.65(1.0) 67.80(1.04) 69.25(2.09) 65.43(1.07) 68.74(1.34) 70.11(0.86)
15% 54.09(2.0) 63.75(1.6) 66.29(0.8) 67.52(1.55) 66.77(2.40) 62.43(1.74) 67.71(1.33) 69.58(1.54)
20% 51.24(1.9) 62.55(1.5) 64.24(1.3) 65.68(1.65) 64.99(2.06) 60.52(1.66) 66.62(1.36) 67.90(1.12)
Cora-ML 5% 70.38(2.1) 77.91(0.9) 78.62(0.8) 79.16(0.99) 79.27(0.64) 77.53(0.70) 79.36(0.98) 80.02(0.82)
10% 68.75(1.7) 76.40(1.1) 76.29(0.9) 77.34(1.01) 77.87(0.66) 76.18(0.88) 77.96(0.82) 78.68(0.90)
15% 67.38(1.1) 75.46(1.4) 74.61(0.6) 76.54(0.83) 76.97(0.90) 74.09(1.15) 77.10(0.56) 77.34(0.92)
20% 65.09(2.0) 73.42(1.7) 73.40(0.8) 75.36(0.80) 75.36(1.28) 72.54(0.89) 75.99(0.52) 76.72(0.79)
Photo 5% 80.79(1.4) 75.80(2.8) 84.86(0.7) 81.91(0.8) 82.28(1.0) 82.25(1.1) 84.18(0.8) 87.09(0.6)
10% 79.11(1.0) 72.74(3.4) 82.27(0.9) 77.49(1.0) 78.12(0.9) 79.17(0.8) 82.21(0.7) 85.04(0.5)
15% 77.83(0.9) 68.99(0.5) 80.39(0.7) 75.28(1.0) 76.98(0.6) 76.68(0.8) 80.78(0.5) 82.11(1.1)
20% 76.97(0.8) 68.59(2.0) 78.88(0.8) 71.85(1.2) 74.92(1.3) 73.10(0.7) 78.90(0.9) 79.21(2.2)
(a) Cora
(b) CiteSeer
(c) Cora-ML
(d) Photo
Figure 4: Performances on graph clustering.
VII-A3 Baseline Models

### v3

@@ V-C Unsupervised Tuning Strategy
[…]
@@
$\small\begin{split}\mathcal{L}_{pnc}^{(t)}=\operatorname{Tr}((\sigma(\mathbf{H}_{1}^{(t)})^{\top}\mathbf{L}^{(t)}\sigma(\mathbf{H}_{1}^{(t)}))\oslash(\sigma(\mathbf{H}_{1}^{(t)})^{\top}\mathbf{D}^{(t)}\sigma(\mathbf{H}_{1}^{(t)})),\end{split}$
where $\sigma(\cdot)$ is the sigmoid function. We record the value of $\mathcal{L}_{pnc}^{(t)}$ for each iteration during training and select the model with the minimum $\mathcal{L}_{pnc}$ as our final result. We denote the best metric as $\mathcal{L}_{pnc}^{t^{*}}$. It is worth noting that the computing of the pseudo normalized cut $\mathcal{L}_{pnc}$ does not rely on the labels, which is suitable for the unsupervised training of GCL.
Figure 5: $\mathcal{L}_{pnc}^{t^{*}}$ v.s. Acc.
Besides, we do not introduce additional parameters to ensure the time efficiency of computing $\mathcal{L}_{pnc}$.
@@
In order to verify the effectiveness of the crafted unsupervised tuning strategy, we present the values of $\mathcal{L}_{pnc}^{t^{*}}$ and node classification accuracy with varying $\eta$ in Fig. 5. The results indicate that in most cases $\mathcal{L}_{pnc}^{t^{*}}$ is negatively related to the node classification accuracy, especially when the value of $\mathcal{L}_{pnc}^{t^{*}}$ falls into the neighborhood of the minimum value, i.e., $\eta\in[0.1,3]$ (the Pearson correlation coefficients $r(\mathcal{L}_{pnc}^{t^{*}},accuracy)=-0.89$). This phenomenon demonstrates that it is reasonable to tune $\eta$ via searching for the best $\mathcal{L}_{pnc}^{t^{*}}$, since the GCL model that achieves the minimum $\mathcal{L}_{pnc}^{t^{*}}$ is likely to get the best accuracy.
VI Experiments
@@
In this section, we introduce the experimental settings and conduct comprehensive experiments to validate the effectiveness and efficiency of our robust model. Specifically, we outline the questions that we seek to answer below:
@@
  • RQ1: Comparing with baseline models, does our model achieve better robust performances?
@@
VI-A Dataset & Settings
TABLE I: Dataset statistics.
@@ VI-A Dataset & Settings
Cora $2708$ $5278$ 77 $1433$
Citeseer $3327$ $4552$ 66 $3703$
Cora-ML $2995$ $8416$ 77 $2879$
Photo $7650$ $119081$ 88 $745$
@@
We use six popular benchmark graph datasets: Cora, CiteSeer, Cora-ML, Photo, Computers and WikiCS [15, 55, 56] for evaluation. To validate the quality of the node embeddings generated by GCL effectively, we train a multi-class logistic regression model with Adam optimizer [57] and randomly split the node labels into three groups: training, validation and testing set with the ratio as $\%, 1010\%$ and $80\%$. We run all the models $10$ times with different seeds and report the mean value for a fair comparison. The vital hyperparameter $\eta$ is tuned based on the unsupervised tuning strategy without querying the node labels.
TABLE II: Performances on node classification against Mettack.
@@
Dataset $\frac{B}{|E|}$ PiGCL BGRL DGI MVGRL GRACE ARIEL GCA SPAN SPAGCL GCIR
Cora 5% 75.25 (0.9) 66.97(1.4) 73.41(1.0) 75.73(0.5) 72.46(1.8) 74.92(1.2) 73.80(1.1) 74.18(0.9) 76.26 (0.8) 78.78(1.0)
10% 70.27 (1.3) 60.11(1.7) 68.03(1.8) 71.10(1.4) 65.37(2.8) 71.93(1.2) 70.26(1.3) 68.81(1.9) 71.33 (0.5) 75.93(0.9)
15% 65.94 (0.6) 51.85(3.0) 59.70(2.7) 64.69(1.9) 52.62(3.1) 52.91(4.4) 57.89(1.5) 58.44(3.2) 63.83 (0.7) 71.89(1.6)
20% 56.54 (1.8) 44.18(4.3) 51.43(2.3) 57.29(2.8) 41.17(4.0) 43.71(3.9) 46.03(3.0) 45.71(2.5) 50.96 (1.0) 68.81(2.4)
CiteSeer 5% 69.60 (1.3) 56.90(2.7) 66.43(1.1) 68.93(0.9) 69.55(2.0) 70.79(1.7) 67.93(1.1) 69.08(1.3) 71.21 (1.2) 70.93(1.3)
10% 65.40 (0.9) 52.30(2.6) 60.78(2.8) 65.01(1.2) 63.49(2.3) 69.14(2.1) 60.82(1.3) 65.18(2.6) 69.60 (1.4) 69.43(1.3)
15% 59.42 (2.1) 44.44(2.7) 53.30(2.4) 58.51(1.3) 53.67(2.7) 56.46(2.3) 52.52(1.7) 56.08(3.5) 62.44 (0.9) 65.84(1.2)
20% 51.78 (0.8) 40.89(3.2) 49.54(1.9) 51.69(1.6) 47.62(3.0) 47.93(3.2) 48.79(1.4) 50.68(3.0) 54.56 (0.6) 62.65(1.5)
Cora-ML 5% 68.59 (0.8) 64.26(1.9) 67.06(2.9) 70.18(2.5) 68.65(2.0) 71.21(1.7) 69.84(1.7) 69.34(0.9) 72.33 (1.1) 77.90(0.8)
10% 49.78 (1.8) 53.08(2.6) 49.34(2.1) 54.20(2.4) 50.49(1.7) 51.87(1.8) 46.81(1.9) 50.44(1.2) 55.60 (0.9) 71.57(1.7)
15% 39.19 (1.3) 44.09(3.6) 42.21(1.5) 45.96(3.3) 40.89(3.4) 41.11(2.3) 37.17(3.6) 41.77(2.4) 47.24 (1.3) 64.56(1.8)
20% 29.45 (0.89) 37.41(3.3) 34.35(0.4) 37.40(2.3) 31.57(3.3) 28.52(2.2) 26.69(2.1) 33.46(2.5) 34.79 (1.3) 56.43(2.5)
Photo 5% 70.12 (1.7) 66.50(2.8) 62.24(3.4) 73.14(1.3) 70.87(1.1) 69.85(2.4) 76.45(2.1) 72.46(2.3) 70.07 (2.0) 78.37(2.2)
10% 50.75 (2.3) 52.62(3.9) 47.76(3.7) 58.63(3.3) 51.39(1.5) 53.92(1.8) 60.57(2.0) 63.46(0.4) 56.98 (1.5) 67.80(2.4)
15% 43.70 (1.) 43.56(2.4) 40.73(3.6) 49.82(2.0) 44.37(1.6) 45.98(1.8) 51.66(1.4) 50.02(0.4) 49.87 (0.9) 59.19(1.5)
20% 27.35 (1.5) 35.68(3.5) 37.91(4.0) 45.00(2.0) 32.45(1.4) 34.19(1.8) 45.35(1.3) 46.81(3.4) 35.03 (1.3) 54.68(2.9)
Computers 5% 62.78 (1.5) 63.12(1.3) 63.91(1.9) 69.16(0.6) 66.22(1.0) 67.51(1.7) 68.02(0.9) 64.06(1.4) 64.85 (1.1) 72.25(2.2)
10% 59.70 (2.5) 61.33(2.1) 57.06(1.9) 64.58(0.4) 58.07(2.0) 57.04(1.5) 58.90(2.1) 54.81(1.3) 57.68 (1.1) 66.14(2.1)
15% 54.26 (0.7) 52.59(3.4) 51.40(2.0) 54.99(2.6) 52.58(2.1) 43.65(2.6) 52.90(2.6) 45.63(3.2) 54.82 (0.8) 59.93(3.8)
20% 49.89 (1.3) 37.61(2.3) 41.54(4.9) 44.98(3.5) 37.21(3.9) 34.71(1.7) 38.81(3.6) 37.39(2.1) 42.87 (0.6) 51.13(4.1)
WikiCS 5% 41.58 (3.12) 40.91(1.0) 43.71(2.2) 42.17(1.8) 41.30(1.1) 43.92(1.0) 41.67(0.8) 40.60(1.1) 43.70 (1.3) 55.93(0.8)
10% 32.27 (1.87) 29.93(0.9) 33.88(2.7) 30.73(0.7) 30.89(1.1) 33.55(1.4) 31.18(0.7) 28.48(0.7) 30.56 (1.2) 49.30(0.6)
15% 25.68 (2.2) 24.35(0.7) 27.10(2.0) 24.54(1.1) 25.40(0.4) 27.37(0.6) 25.25(0.5) 23.32(0.7) 27.32 (1.5) 42.96(1.1)
20% 23.83 (1.84) 21.60(1.8) 23.79(1.3) 22.38(0.8) 23.23(0.4) 24.56(0.7) 23.03(0.7) 21.79(0.5) 23.94 (0.9) 36.47(1.0)
TABLE III: Performances on node classification against CLGA.
@@
Dataset $\frac{B}{|E|}$ PiGCL BGRL DGI MVGRL GRACE ARIEL GCA SPAN SPAGCL GCIR
Cora 5% 79.72 (0.9) 70.78(1.1) 78.96(1.3) 80.09(0.9) 78.95(0.9) 78.47(0.7) 78.55(0.7) 78.96(0.7) 77.46 (1.0) 79.13(0.6)
10% 78.11 (1.2) 68.35(1.9) 75.64(1.2) 77.92(0.9) 76.79(1.4) 76.31(0.9) 77.15(0.8) 77.22(1.1) 75.93 (0.9) 78.62(1.1))
15% 77.01 (2.3) 66.15(1.3) 73.72(0.7) 75.89(1.0) 75.38(1.1) 74.75(0.9) 74.71(1.2) 76.11(1.2) 74.65 (1.6) 78.18(1.0)
20% 75.20 (1.8) 64.44(1.8) 71.94(1.7) 74.64(0.8) 73.13(1.1) 74.44(1.2) 71.78(1.0) 74.43(1.6) 73.74 (1.4) 76.82(1.2)
CiteSeer 5% 68.95 (2.3) 58.52(1.3) 66.24(1.5) 68.60(1.2) 69.38(1.2) 70.43(1.2) 67.92(1.0) 70.38(1.1) 70.14 (1.4) 71.18(1.2)
10% 68.48 (1.8) 56.27(1.3) 65.62(1.5) 67.65(1.0) 67.80(1.0) 69.25(2.1) 65.43(1.1) 68.74(1.3) 68.36 (0.9) 70.11(0.9)
15% 67.30 (1.5) 54.09(2.0) 63.75(1.6) 66.29(0.8) 67.52(1.6) 66.77(2.4) 62.43(1.7) 67.71(1.3) 67.18 (1.0) 69.58(1.5)
20% 64.10 (1.8) 51.24(1.9) 62.55(1.5) 64.24(1.3) 65.68(1.7) 64.99(2.1) 60.52(1.7) 66.62(1.4) 66.58 (1.1) 67.90(1.1)
Cora-ML 5% 78.38 (1.2) 70.38(2.1) 77.91(0.9) 78.62(0.8) 79.16(1.0) 79.27(0.6) 77.53(0.7) 79.36(1.0) 79.18 (1.3) 80.02(0.8)
10% 77.89 (0.7) 68.75(1.7) 76.40(1.1) 76.29(0.9) 77.34(1.0) 77.87(0.7) 76.18(0.9) 77.96(0.8) 77.62 (1.3) 78.68(0.9)
15% 76.77 (1.6) 67.38(1.1) 75.46(1.4) 74.61(0.6) 76.54(0.8) 76.97(0.9) 74.09(1.2) 77.10(0.6) 76.38 (0.8) 77.34(0.9)
20% 75.71 (1.2) 65.09(2.0) 73.42(1.7) 73.40(0.8) 75.36(0.8) 75.36(1.3) 72.54(0.9) 75.99(0.5) 75.22 (1.5) 76.72(0.8)
Photo 5% 77.78 (1.2) 80.79(1.4) 75.80(2.8) 84.86(0.7) 81.91(0.8) 82.28(1.0) 82.25(1.1) 84.18(0.8) 83.62 (0.6) 87.09(0.6)
10% 75.33 (0.9) 79.11(1.0) 72.74(3.4) 82.27(0.9) 77.49(1.0) 78.12(0.9) 79.17(0.8) 82.21(0.7) 81.86 (0.9) 85.04(0.5)
15% 70.97 (1.8) 77.83(0.9) 68.99(0.5) 80.39(0.7) 75.28(1.0) 76.98(0.6) 76.68(0.8) 80.78(0.5) 80.83 (1.1) 82.11(1.1)
20% 67.15 (1.6) 76.97(0.8) 68.59(2.0) 78.88(0.8) 71.85(1.2) 74.92(1.3) 73.10(0.7) 78.90(0.9) 77.66 (1.1) 79.21(2.2)
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
Figure 6: Sensitivity analysis on $\eta$.
VI-B Baselines